# A Two-Timescale Primal-Dual Framework for Reinforcement Learning via Online Dual Variable Guidance

### v2

@@ Appendix C Asynchronous Convergence Rate: Proof Details
Proof of Proposition 4.4
The stated contraction follows from [Zeng et al., 2024, Proposition 1] with two modifications. First, we work on the high-probability visitation event $\mathcal{G}_{\delta}$ and use Lemma C.1 to control the Markovian bias. Second, we adapt the argument to maximization of the reduced objective $f$. By Lemma C.5, $f$ is $\mu_{\mathrm{opt}}$-strongly concave and thus satisfies a Polyak-Łojasiewicz inequality of the form $f(\rho^{\star})-f(\rho)\leq\tfrac{1}{2\mu_{\mathrm{opt}}}\|\nabla f(\rho)\|_{2}^{2}$ for all $\rho\in H$. Substituting this PŁ-inequality into the standard smoothness-based descent argument (with the sign flipped for maximization) and tracking the critic error and mixing terms as in [Zeng et al., 2024, Proposition 1] yields the stated recursion. $\blacksquare$
@@ C.3 Auxiliary Results and Proofs
This section contains additional auxiliary results required for the convergence rate proof.
Lemma C.6 (Direct bound on $\tau_{\mathrm{inh}}(\zeta)$ and $\tau_{\min}$).
Let Assumption 3 hold. Let $\tau_{\mathrm{min}}$ be as defined in (41a) and let $\tau_{\mathrm{inh}}(\zeta)$ be the inhomogeneous mixing time of $\{\mathcal{P}_{\rho}\}$ as defined in (41b). It holds that there exists $M\in\mathbb{N}$ and $\epsilon_{0}\in(0,1)$, such that for any $\zeta\in(0,1)$
$\tau_{\mathrm{inh}}(\zeta)\ \leq\ (M-1)\ +\ M\Big\lceil\frac{\log(1/\zeta)}{\log\!\big(1/(1-\epsilon)\big)}\Big\rceil,\qquad\epsilon=|\mathcal{A}|^{M}(\pi_{\min})^{M}\epsilon_{0},$
and $\tau_{\min}:=\inf_{0\leq\zeta<1}\tau_{\mathrm{inh}}(\zeta)\Big(\tfrac{2-\zeta}{1-\zeta}\Big)^{2}\leq 9\big[(M-1)+M\lceil\log 2/((\pi_{\min})^{M}\epsilon_{0})\rceil\big]$.
Proof  Let Assumption 3 hold and $M\in\mathbb{N}$, $\epsilon_{0}\in(0,1)$ and $\upsilon$ be the positive probability measure on the state space as derived in the proof of Lemma 4.1. Recall the notations $\tilde{\mathcal{P}}_{\rho}$ for the state-transition kernel induced by $\pi_{\rho}$ and $\mathcal{P}_{d}$ for the state-transition kernel induced by the deterministic ergodic policy $\pi_{d}$, that is,
$\tilde{\mathcal{P}}_{\rho}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\pi_{\rho}(a|s)\mathcal{P}(s^{\prime}|s,a),\qquad\mathcal{P}_{d}(s^{\prime}|s)=\mathcal{P}\big(s^{\prime}|s,\pi_{d}(s)\big).$
Arguments identical to the proof of Lemma 4.1 also yield a uniform $M$-step Doeblin minorization for the inhomogeneous state chain. Consider the block consisting of the iterations $i,\dots,i+M-1$. Then, for all $s\in\mathcal{S}$
$\prod_{j=i}^{i+M-1}\tilde{\mathcal{P}}_{\rho_{j}}(\cdot|s)\ \geq\ |\mathcal{A}|^{M}\big(\pi_{\min}\big)^{M}\,\mathcal{P}_{d}^{\,M}(\cdot|s)\ \geq\ \epsilon\,\upsilon(\cdot),\qquad\epsilon:=|\mathcal{A}|^{M}(\pi_{\min})^{M}\epsilon_{0},$
where the last step is the application of the $M$-step Doeblin minorization. By the non-expansiveness of the total variation distance, the Dobrushin coefficients also hold for the induced state-action chain, and consequently, each length-$M$ block has a Dobrushin coefficient at most $1-\epsilon$, and for any $k=mM+r$ with $0\leq r<M$,
$\sup_{x,y\in\mathcal{X}}\Big\|\delta_{x}\prod_{j=i}^{i+k-1}\mathcal{P}_{\rho_{j}}-\delta_{y}\prod_{j=i}^{i+k-1}\mathcal{P}_{\rho_{j}}\Big\|_{\mathrm{TV}}\ \leq\ (1-\epsilon)^{m}.$
Imposing $(1-\epsilon)^{m}\leq\zeta$ gives
$\tau_{\mathrm{inh}}(\zeta)\ \leq\ (M-1)\ +\ M\Big\lceil\frac{\log(1/\zeta)}{\log\!\big(1/(1-\epsilon)\big)}\Big\rceil,\qquad\epsilon=|\mathcal{A}|^{M}(\pi_{\min})^{M}\epsilon_{0},$
@@
$\tau_{\min}\ :=\ \inf_{0\leq\zeta<1}\ \tau_{\mathrm{inh}}(\zeta)\Big(\tfrac{2-\zeta}{1-\zeta}\Big)^{2}\ \leq\ 9\cdot\tau_{\mathrm{inh}}(1/2)\ \leq\ 9\Big[(M-1)\ +\ M\Big\lceil\tfrac{\log 2}{\log\!\big(1/(1-\epsilon)\big)}\Big\rceil\Big].$
Using $\log(1-\epsilon)\geq\epsilon$ further yields the simpler bound
$\tau_{\min}\ \leq\ 9\big[(M-1)+M\lceil\log 2/(|\mathcal{A}|^{M}(\pi_{\min})^{M}\epsilon_{0})\rceil\big].$
$\blacksquare$
@@
The Lipschitz property of the dual-induced transition kernel has been established in the proof of Lemma B.3. We next provide a proof of the Lipschitz continuity of the dual-induced stationary distribution which is applied in the proof of the visitation floor lemma.
Lemma C.7 (Lipschitz stationary law under UGE).
@@
Proof.
Under Assumption 3, for each $\rho\in H$ there is a unique stationary distribution $\mu_{\rho}$ and applying [Mitrophanov, 2005, Corollary 3.1] we have
@@
where $L_{\mu}:=\log_{\varrho}(C_{\mathcal{X}}^{-1})+\frac{1}{1-\varrho}$. It remains to control $\|\mathcal{P}_{\rho}-\mathcal{P}_{\hat{\rho}}\|_{\infty}$ in terms of $\rho-\hat{\rho}$, which has already been shown in the proof of Lemma B.3, where the Lipschitz property of the $\mathcal{P}_{\rho}$ has been established. This shows (57). The perturbation bound (58) follows from the triangle inequality applied as
@@
Applying Lemma B.3 gives the bound on the policy difference by $C_{\mathcal{Q}}$ times the $\rho$-difference per state. ∎
The following lemma provides the bound on the term $\mathbb{E}[\langle V_{k}-\lambda(\rho_{k}),\Delta^{X}_{k+1}\rangle|\mathcal{G}_{\delta}]$ that appears in step (III.) of the proof of Proposition 4.3. Recall that $\Delta^{X}_{k+1}$, defined in (55), denotes the mismatch between the law driving the update index selection and the ideal, mixed law.
Lemma C.8.
Let $K(\delta),\,\mathcal{G}_{\delta}$ be as in Prop. 4.2 and $\mathcal{K}$ as in (24). Moreover, let the Lipschitz envelope $L$ and the gradient bounds B,DB,\,D be as in Prop. 4.3. For any $k\geq\max\{K(\delta),\mathcal{K}\}$, it holds that
@@
Proof
The Lemma is an adjusted version of [Zeng et al., 2024, Lemma 2], with the difference lying in the boundedness of the gradient in our projected setting. The proof is also closely related to theirs, with changes occurring at the points where a uniform bound on the gradient can be applied.
The idea for handling the Markovian mismatch is to compare the inhomogeneous index process $\{X_{t}\}$ to an auxiliary chain frozen at parameter $\rho_{k-\tau_{k}}$, (see also the proof of [Zeng et al., 2024, Lemma 1]). Specifically, define

### v3

@@ C.2 Proofs of the Propositions 18 and 19
Proof of Proposition 19
Fix $\delta\in(0,1)$ and let $k\geq\max\{K(\delta),\mathcal{K}\}$. We work on the high-probability event $\mathcal{G}_{\delta}$. Since $\tau_{k}\geq 1$, the burn-in condition $\tau_{k}\alpha_{\mathrm{env},k-\tau_{k}}\leq\tfrac{1}{\sqrt{3}LB}$ implies $\alpha_{\mathrm{env},k-\tau_{k}}\leq\tfrac{1}{\sqrt{3}LB}.$ For the chosen two-timescale stepsizes, $\beta_{t}\leq\alpha_{t}$ for all sufficiently large $t$, hence
$\bar{\beta}_{k}\leq\beta_{\mathrm{env},k}\leq\beta_{\mathrm{env},k-\tau_{k}}\leq\alpha_{\mathrm{env},k-\tau_{k}}\leq\frac{1}{\sqrt{3}LB}\leq\frac{1}{L},$
where the last inequality follows from $B\geq 1$. Similarly to the proof of Proposition 18, we set
$\widehat{H}_{k}:=\widehat{h}_{k}(\rho_{k-1},V_{k-1};X_{k}),\qquad\widetilde{H}_{k}:=\mathbb{E}[\widehat{H}_{k}\mid\mathcal{F}_{k-1}].$
Then, by taking the expectation under the stationary distribution in (17b), we have
$\mathbb{E}_{X\sim\mu_{\rho_{k-1}}}\Big[\widehat{h}_{k}(\rho_{k-1},V_{k-1};X)\Big]=\operatorname{diag}(\mu_{\rho_{k-1}})\nabla_{\rho}L(V_{k-1},\rho_{k-1}).$
Using this stationary mean-field identity and $\nabla f(\rho)=\nabla_{\rho}L(\lambda(\rho),\rho)$, we decompose
$\widetilde{H}_{k+1}=D(\rho_{k})\nabla f(\rho_{k})+\Delta^{V}_{k+1}+\bar{\Delta}^{X}_{k+1},$
where
$D(\rho_{k}):=\operatorname{diag}(\mu_{\rho_{k}}),\quad\Delta^{V}_{k+1}:=\mathbb{E}_{X\sim\mu_{\rho_{k}}}\Big[\widehat{h}_{k+1}(\rho_{k},V_{k};X)-\widehat{h}_{k+1}(\rho_{k},\lambda(\rho_{k});X)\Big],$
and
$\bar{\Delta}^{X}_{k+1}:=\widetilde{H}_{k+1}-\mathbb{E}_{X\sim\mu_{\rho_{k}}}\big[\widehat{h}_{k+1}(\rho_{k},\lambda(\rho_{k});X)\big].$
To isolate the core finite-time contraction mechanism, we carry out the following one-step estimate in the regime where the clipping is inactive after burn-in. By $L$-smoothness of $f$,
$f(\rho_{k+1})$ $\geq f(\rho_{k})+\bar{\beta}_{k+1}\langle\nabla f(\rho_{k}),\widehat{H}_{k+1}\rangle-\frac{L}{2}\bar{\beta}_{k+1}^{2}\|\widehat{H}_{k+1}\|_{2}^{2}.$
By Proposition 15(iv), $\|\widehat{H}_{k+1}\|^{2}_{2}\leq B^{2}$. Substituting the decomposition of $\widetilde{H}_{k+1}$, we obtain
$f(\rho_{k+1})$ $\geq f(\rho_{k})+\bar{\beta}_{k+1}\langle\nabla f(\rho_{k}),D(\rho_{k})\nabla f(\rho_{k})\rangle$
$\qquad+\bar{\beta}_{k+1}\langle\nabla f(\rho_{k}),\Delta^{V}_{k+1}+\bar{\Delta}^{X}_{k+1}+(\widehat{H}_{k+1}-\widetilde{H}_{k+1})\rangle+\frac{LB^{2}}{2}\bar{\beta}_{k+1}^{2}.$
Rearranging gives
$f(\rho^{\star})-f(\rho_{k+1})$ $\leq f(\rho^{\star})-f(\rho_{k})-\bar{\beta}_{k+1}\langle\nabla f(\rho_{k}),D(\rho_{k})\nabla f(\rho_{k})\rangle$
$\qquad-\bar{\beta}_{k+1}\langle\nabla f(\rho_{k}),\Delta^{V}_{k+1}+\bar{\Delta}^{X}_{k+1}+(\widehat{H}_{k+1}-\widetilde{H}_{k+1})\rangle+\frac{LB^{2}}{2}\bar{\beta}_{k+1}^{2}.$ (57)
Using $\mu_{\rho_{k}}(s,a)\geq p_{\star}$ for all $(s,a)\in\mathcal{X}$
$-\langle\nabla f(\rho_{k}),D(\rho_{k})\nabla f(\rho_{k})\rangle$ $\leq-p_{\star}\|\nabla f(\rho_{k})\|_{2}^{2}\leq-2p_{\star}\mu_{\mathrm{opt}}\bigl(f(\rho^{\star})-f(\rho_{k})\bigr).$ (58)
By Lipschitz continuity of $\widehat{h}_{k}(\rho,\cdot;X)$ in $V$, $\|\Delta^{V}_{k+1}\|_{2}\leq L\|V_{k}-\lambda(\rho_{k})\|_{2}.$ Using Young’s inequality with parameter $p_{\star}$,
$\langle\nabla f(\rho_{k}),\Delta^{V}_{k+1}\rangle\leq\frac{p_{\star}}{2}\|\nabla f(\rho_{k})\|_{2}^{2}+\frac{1}{2p_{\star}}\|\Delta^{V}_{k+1}\|_{2}^{2},$
and therefore
$-\bar{\beta}_{k+1}\langle\nabla f(\rho_{k}),\Delta^{V}_{k+1}\rangle\leq\frac{p_{\star}}{2}\bar{\beta}_{k+1}\|\nabla f(\rho_{k})\|_{2}^{2}+\frac{L^{2}}{2p_{\star}}\bar{\beta}_{k+1}\|V_{k}-\lambda(\rho_{k})\|_{2}^{2}.$ (59)
By the Polyak–Łojasiewicz inequality for $\mu_{\mathrm{opt}}$-strongly concave $f$,
$\tfrac{1}{2}\|\nabla f(\rho)\|_{2}^{2}\geq\mu_{\mathrm{opt}}\bigl(f(\rho^{\star})-f(\rho)\bigr).$
Combining (59) with (58) and the Polyak–Łojasiewicz inequality and inserting the bounds into Equation (57) gives
$f(\rho^{\star})-f(\rho_{k+1})$ $\leq(1-p_{\star}\mu_{\mathrm{opt}}\bar{\beta}_{k+1})\bigl(f(\rho^{\star})-f(\rho_{k})\bigr)+\frac{L^{2}}{2p_{\star}}\bar{\beta}_{k+1}\|V_{k}-\lambda(\rho_{k})\|_{2}^{2}$
$\qquad-\bar{\beta}_{k+1}\langle\nabla f(\rho_{k}),\bar{\Delta}^{X}_{k+1}+(\widehat{H}_{k+1}-\widetilde{H}_{k+1})\rangle+\frac{LB^{2}}{2}\bar{\beta}_{k+1}^{2}.$ (60)
To address the Martingale difference term $\widehat{H}_{k+1}-\widetilde{H}_{k+1}$, note that
$\mathbb{E}[\langle\nabla f(\rho_{k}),\widehat{H}_{k+1}-\widetilde{H}_{k+1}\rangle|\mathcal{F}_{k}]=0$
Taking expectation conditional on $\mathcal{G}_{\delta}$, using $\beta_{k}\leq\bar{\beta}_{k+1}\leq\beta_{\mathrm{env},k}$, and invoking the Markovian mismatch bound from Lemma 34,
$\mathbb{E}\!\left[\langle\nabla f(\rho_{k}),\bar{\Delta}^{X}_{k+1}\rangle|\mathcal{G}_{\delta}\right]\leq 12L^{2}B^{3}\tau_{k}^{2}\beta_{\mathrm{env},k-\tau_{k}},$
we arrive at
$\mathbb{E}[f(\rho^{\star})-f(\rho_{k+1})|\mathcal{G}_{\delta}]$ $\leq\bigl(1-p_{\star}\mu_{\mathrm{opt}}\beta_{k}\bigr)\mathbb{E}[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}]+\frac{L^{2}\beta_{\mathrm{env},k}}{2p_{\star}}\mathbb{E}[\|V_{k}-\lambda(\rho_{k})\|_{2}^{2}|\mathcal{G}_{\delta}]$
$\qquad+12L^{2}B^{3}\tau_{k}^{2}\beta_{\mathrm{env},k-\tau_{k}}\beta_{\mathrm{env},k}+\frac{L}{2}B^{2}\beta_{\mathrm{env},k}^{2}.$
The tower property with $\mathcal{G}_{\delta}$ eliminates the Martingale difference term. Finally, since $\tau_{k}\geq 1$ and $\beta_{\mathrm{env},k-\tau_{k}}\geq\beta_{\mathrm{env},k}$, $\beta_{\mathrm{env},k}^{2}\leq\tau_{k}^{2}\beta_{\mathrm{env},k-\tau_{k}}\beta_{\mathrm{env},k},$ so the last term can be absorbed into the same remainder term. Thus,
$\mathbb{E}[f(\rho^{\star})-f(\rho_{k+1})|\mathcal{G}_{\delta}]$ $\leq\bigl(1-p_{\star}\mu_{\mathrm{opt}}\beta_{k}\bigr)\mathbb{E}[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}]+\frac{L^{2}}{2p_{\star}}\beta_{\mathrm{env},k}\mathbb{E}[\|V_{k}-\lambda(\rho_{k})\|_{2}^{2}|\mathcal{G}_{\delta}]$
$\qquad+\frac{25L^{2}B^{3}}{2}\tau_{k}^{2}\beta_{\mathrm{env},k-\tau_{k}}\beta_{\mathrm{env},k},$
which is the claimed recursion. $\blacksquare$
@@ C.3 Auxiliary Results and Proofs
This section contains additional auxiliary results required for the convergence rate proof as well as the proof of Corollary 17.
We first show the Lipschitz-continuity of the mapping $\pi\mapsto V^{\pi}_{ur}$.
Lemma 30 (Lipschitz continuity of the unregularized value in the policy)
For
a stationary policy $\pi\in\Pi$, let $V^{\pi}_{ur}$ denote the unregularized discounted value function. Then, for any two stationary policies $\pi,\hat{\pi}\in\Pi$,
$\|V^{\pi}_{ur}-V^{\hat{\pi}}_{ur}\|_{\infty}\leq\frac{C_{r}}{(1-\gamma)^{2}}\max_{s\in\mathcal{S}}\|\pi(\cdot|s)-\hat{\pi}(\cdot|s)\|_{1}.$
Proof  For a stationary policy $\pi$, define
$r_{\pi}(s):=\sum_{a\in\mathcal{A}}\pi(a|s)r(s,a),\qquad[\mathcal{P}_{\pi}]_{ss^{\prime}}:=\sum_{a\in\mathcal{A}}\pi(a|s)\mathcal{P}(s^{\prime}|s,a).$
Then, $V^{\pi}_{ur}$ satisfies the Bellman equation
$V^{\pi}_{ur}=r_{\pi}+\gamma\mathcal{P}_{\pi}V^{\pi}_{ur}.$
Subtracting the Bellman equations for $\pi$ and $\hat{\pi}$ yields
$V^{\pi}_{ur}-V^{\hat{\pi}}_{ur}=\gamma\mathcal{P}_{\pi}\bigl(V^{\pi}_{ur}-V^{\hat{\pi}}_{ur}\bigr)+(r_{\pi}-r_{\hat{\pi}})+\gamma(\mathcal{P}_{\pi}-\mathcal{P}_{\hat{\pi}})V^{\hat{\pi}}_{ur}$.
Hence,
$V^{\pi}_{ur}-V^{\hat{\pi}}_{ur}=(I-\gamma\mathcal{P}_{\pi})^{-1}\Bigl[(r_{\pi}-r_{\hat{\pi}})+\gamma(\mathcal{P}_{\pi}-\mathcal{P}_{\hat{\pi}})V^{\hat{\pi}}_{ur}\Bigr$].
Since $\mathcal{P}_{\pi}$ is a stochastic matrix, we have $\|\mathcal{P}_{\pi}^{t}v\|_{\infty}\leq\|v\|_{\infty}$ for all $t\geq 0$ and $v\in\mathbb{R}^{|\mathcal{S}|}$. Hence, using the Neumann-series expansion,
$(I-\gamma\mathcal{P}_{\pi})^{-1}v=\sum_{t=0}^{\infty}\gamma^{t}\mathcal{P}_{\pi}^{t}v,$
it follows that
$\|(I-\gamma\mathcal{P}_{\pi})^{-1}v\|_{\infty}\leq\sum_{t=0}^{\infty}\gamma^{t}\|\mathcal{P}_{\pi}^{t}v\|_{\infty}\leq\sum_{t=0}^{\infty}\gamma^{t}\|v\|_{\infty}=\frac{1}{1-\gamma}\|v\|_{\infty},\qquad\forall\,v\in\mathbb{R}^{|\mathcal{S}|}.$
See also Puterman (1994, Theorem 6.1.1) for the corresponding discounted-value representation. Moreover, for every $s\in\mathcal{S}$,
$|r_{\pi}(s)-r_{\hat{\pi}}(s)|=\left|\sum_{a\in\mathcal{A}}(\pi(a|s)-\hat{\pi}(a|s))r(s,a)\right|\leq C_{r}\|\pi(\cdot|s)-\hat{\pi}(\cdot|s)\|_{1},$
@@
$\|r_{\pi}-r_{\hat{\pi}}\|_{\infty}\leq C_{r}\max_{s\in\mathcal{S}}\|\pi(\cdot|s)-\hat{\pi}(\cdot|s)\|_{1}.$
Likewise, for any $v\in\mathbb{R}^{|\mathcal{S}|}$ and $s\in\mathcal{S}$,
$|((\mathcal{P}_{\pi}-\mathcal{P}_{\hat{\pi}})v)(s)|$ $=\left|\sum_{a\in\mathcal{A}}(\pi(a|s)-\hat{\pi}(a|s))\sum_{s^{\prime}\in\mathcal{S}}\mathcal{P}(s^{\prime}|s,a)v(s^{\prime})\right|$
$\leq\|v\|_{\infty}\|\pi(\cdot|s)-\hat{\pi}(\cdot|s)\|_{1}.$
Hence,
$\|(\mathcal{P}_{\pi}-\mathcal{P}_{\hat{\pi}})v\|_{\infty}\leq\|v\|_{\infty}\max_{s\in\mathcal{S}}\|\pi(\cdot|s)-\hat{\pi}(\cdot|s)\|_{1}.$
Since $0\leq r(s,a)\leq C_{r}$, we also have
$\|V^{\hat{\pi}}_{ur}\|_{\infty}\leq\frac{C_{r}}{1-\gamma}.$
Combining the above bounds gives
[…]
$\leq\frac{1}{1-\gamma}\left(C_{r}+\gamma\frac{C_{r}}{1-\gamma}\right)\max_{s\in\mathcal{S}}\|\pi(\cdot|s)-\hat{\pi}(\cdot|s)\|_{1}$
$=\frac{C_{r}}{(1-\gamma)^{2}}\max_{s\in\mathcal{S}}\|\pi(\cdot|s)-\hat{\pi}(\cdot|s)\|_{1},$
which proves the claim. $\blacksquare$
Proof of Corollary 17.
By Proposition 4, the optimal dual solution $\rho^{\star}$ induces the optimal regularized policy, that is,
$\pi_{\rho^{\star}}=\pi_{r}^{\star}.$
By the definition of $L_{\pi}$ in Corollary 17, whose existence follows directly from the Lipschitz estimate derived in the proof of Lemma 23, we have
$\max_{s\in\mathcal{S}}\|\pi_{\rho_{k}}(\cdot|s)-\pi_{\rho^{\star}}(\cdot|s)\|_{1}\leq L_{\pi}\|\rho_{k}-\rho^{\star}\|_{2}.$
Applying Lemma 30 with $\pi=\pi_{\rho_{k}}$ and $\hat{\pi}=\pi_{\rho^{\star}}=\pi_{r}^{\star}$, we obtain
$\|V^{\pi_{\rho_{k}}}_{ur}-V^{\pi_{r}^{\star}}_{ur}\|_{\infty}\leq\frac{C_{r}}{(1-\gamma)^{2}}\max_{s\in\mathcal{S}}\|\pi_{\rho_{k}}(\cdot|s)-\pi_{\rho^{\star}}(\cdot|s)\|_{1}\leq L_{V}\|\rho_{k}-\rho^{\star}\|_{2}.$
Next, by Proposition 1 and the specialization $0\leq G(\pi(\cdot|s))\leq\log|\mathcal{A}|$ for the Shannon entropy regularizer, it holds for every $s\in\mathcal{S}$ that
$0\leq V^{\star}_{ur}(s)-V^{\pi_{r}^{\star}}_{ur}(s)\leq\frac{\eta_{\rho}\log|\mathcal{A}|}{1-\gamma}=B_{\mathrm{reg}}.$
Therefore,
$V^{\star}_{ur}(s)-V^{\pi_{\rho_{k}}}_{ur}(s)$ $=\bigl(V^{\star}_{ur}(s)-V^{\pi_{r}^{\star}}_{ur}(s)\bigr)+\bigl(V^{\pi_{r}^{\star}}_{ur}(s)-V^{\pi_{\rho_{k}}}_{ur}(s)\bigr)$
$\leq B_{\mathrm{reg}}+\|V^{\pi_{r}^{\star}}_{ur}-V^{\pi_{\rho_{k}}}_{ur}\|_{\infty}.$
Since $V^{\star}_{ur}(s)-V^{\pi_{\rho_{k}}}_{ur}(s)\geq 0$, this implies
$0\leq\left(V^{\star}_{ur}(s)-V^{\pi_{\rho_{k}}}_{ur}(s)-B_{\mathrm{reg}}\right)_{+}\leq\|V^{\pi_{r}^{\star}}_{ur}-V^{\pi_{\rho_{k}}}_{ur}\|_{\infty}\leq L_{V}\|\rho_{k}-\rho^{\star}\|_{2}.$
Squaring both sides and taking conditional expectations on $\mathcal{G}_{\delta}$ yields
$\mathbb{E}\!\left[\left(V^{\star}_{ur}(s)-V^{\pi_{\rho_{k}}}_{ur}(s)-B_{\mathrm{reg}}\right)_{+}^{2}\;\middle|\;\mathcal{G}_{\delta}\right]\leq L_{V}^{2}\mathbb{E}\!\left[\|\rho_{k}-\rho^{\star}\|_{2}^{2}\;\middle|\;\mathcal{G}_{\delta}\right].$
The final rate statement follows immediately from Theorem 16. $\blacksquare$
Lemma 31 (Direct bound on $\tau_{\mathrm{inh}}(\zeta)$ and $\tau_{\min}$)
Let Assumption 13(ii) hold. Let $\tau_{\min}$ be as defined in (41a) and let $\tau_{\mathrm{inh}}(\zeta)$ be the inhomogeneous mixing time of $\{P_{\rho_{k}}\}$ as defined in (41b). Then, for any $\zeta\in(0,1)$,
$\tau_{\mathrm{inh}}(\zeta)\leq(m_{\star}-1)+m_{\star}\left\lceil\frac{\log(1/\zeta)}{\log(1/\kappa)}\right\rceil,$
and consequently
$\tau_{\min}=\inf_{0\leq\zeta<1}\tau_{\mathrm{inh}}(\zeta)\Bigl(\frac{2-\zeta}{1-\zeta}\Bigr)^{2}\leq 9\Biggl[(m_{\star}-1)+m_{\star}\left\lceil\frac{\log 2}{\log(1/\kappa)}\right\rceil\Biggr].$
Proof  By Assumption 13(ii), there exist constants $m_{\star}\in\mathbb{N}$ and $\kappa\in(0,1)$ such that for every sequence $\{\rho_{t}\}_{t\in\mathbb{N}}\subset H$,
$\sup_{k\geq 0}\delta\!\left(\prod_{t=k}^{k+m_{\star}-1}P_{\rho_{t}}\right)\leq\kappa.$
Fix an initial time index $i\geq 0$, and let $k=q\,m_{\star}+r$ with $q\in\mathbb{N}_{0}$ and $0\leq r<m_{\star}$. Write the product of $k$ consecutive kernels as
$\prod_{t=i}^{i+k-1}P_{\rho_{t}}=\left(\prod_{b=0}^{q-1}\prod_{t=i+bm_{\star}}^{i+(b+1)m_{\star}-1}P_{\rho_{t}}\right)\left(\prod_{t=i+qm_{\star}}^{i+qm_{\star}+r-1}P_{\rho_{t}}\right).$
Using the submultiplicativity of Dobrushin’s ergodic coefficient,
$\delta(Q_{1}Q_{2})\leq\delta(Q_{1})\delta(Q_{2}),$
together with the bound $\delta(Q)\leq 1$ for every Markov kernel $Q$, we obtain
$\delta\!\left(\prod_{t=i}^{i+k-1}P_{\rho_{t}}\right)$ $\leq\prod_{b=0}^{q-1}\delta\!\left(\prod_{t=i+bm_{\star}}^{i+(b+1)m_{\star}-1}P_{\rho_{t}}\right)\cdot\delta\!\left(\prod_{t=i+qm_{\star}}^{i+qm_{\star}+r-1}P_{\rho_{t}}\right)$
$\leq\kappa^{q}.$
Hence, for all $x,y\in\mathcal{X}$,
$\left\|\delta_{x}\prod_{t=i}^{i+k-1}P_{\rho_{t}}-\delta_{y}\prod_{t=i}^{i+k-1}P_{\rho_{t}}\right\|_{\mathrm{TV}}\leq\delta\!\left(\prod_{t=i}^{i+k-1}P_{\rho_{t}}\right)\leq\kappa^{q}.$
Therefore, if $\kappa^{q}\leq\zeta$, then any product of length $k=q\,m_{\star}+r$ contracts total variation distance by at most $\zeta$. Choosing
$q=\left\lceil\frac{\log(1/\zeta)}{\log(1/\kappa)}\right\rceil$
yields
$\tau_{\mathrm{inh}}(\zeta)\leq(m_{\star}-1)+m_{\star}\left\lceil\frac{\log(1/\zeta)}{\log(1/\kappa)}\right\rceil.$
For $\tau_{\min}$, we proceed exactly as in the current argument and evaluate at $\zeta=\tfrac{1}{2}$:
$\tau_{\min}=\inf_{0\leq\zeta<1}\tau_{\mathrm{inh}}(\zeta)\Bigl(\frac{2-\zeta}{1-\zeta}\Bigr)^{2}\leq\tau_{\mathrm{inh}}(1/2)\Bigl(\frac{2-1/2}{1-1/2}\Bigr)^{2}=9\,\tau_{\mathrm{inh}}(1/2).$
Combining this with the previous bound gives
$\tau_{\min}\leq 9\Biggl[(m_{\star}-1)+m_{\star}\left\lceil\frac{\log 2}{\log(1/\kappa)}\right\rceil\Biggr],$
which proves the claim. $\blacksquare$
@@
The Lipschitz property of the dual-induced transition kernel has been established in the proof of Lemma 23. We next provide a proof of the Lipschitz continuity of the dual-induced stationary distribution which is applied in the proof of the visitation floor lemma.
Lemma 32 (Lipschitz stationary law under Assumption 13)
@@
Proof  Under Assumption 13, for each $\rho\in H$ there is a unique stationary distribution $\mu_{\rho}$ and applying (Mitrophanov, 2005, Corollary 3.1) we have
@@
where $L_{\mu}:=\log_{\varrho}(C_{\mathcal{X}}^{-1})+\frac{1}{1-\varrho}$. It remains to control $\|\mathcal{P}_{\rho}-\mathcal{P}_{\hat{\rho}}\|_{\infty}$ in terms of $\rho-\hat{\rho}$, which has already been shown in the proof of Lemma 23, where the Lipschitz property of the $\mathcal{P}_{\rho}$ has been established. This shows (61). The perturbation bound (62) follows from the triangle inequality applied as
@@
Applying Lemma 23 gives the bound on the policy difference by $C_{\mathcal{Q}}$ times the $\rho$-difference per state. $\blacksquare$
The following lemma provides the bound on the term $\mathbb{E}[\langle V_{k}-\lambda(\rho_{k}),\Delta^{X}_{k+1}\rangle|\mathcal{G}_{\delta}]$ that appears in step (III.) of the proof of Proposition 18. Recall that $\Delta^{X}_{k+1}$, defined in (55), denotes the mismatch between the law driving the update index selection and the ideal, mixed law.
Lemma 33
Let $K(\delta),\,\mathcal{G}_{\delta}$ be as in Prop. 15 and $\mathcal{K}$ as in (24). Moreover, let the Lipschitz envelope $L$ and the gradient bounds B,DB,\,D be as in Prop. 18. For any $k\geq\max\{K(\delta),\mathcal{K}\}$, it holds that
@@
Proof  The lemma is an adjusted version of (Zeng et al., 2024, Lemma 2), with the difference lying in the boundedness of the gradient in our projected setting. The proof is also closely related to theirs, with changes occurring at the points where a uniform bound on the gradient can be applied.
The idea for handling the Markovian mismatch is to compare the inhomogeneous index process $\{X_{t}\}$ to an auxiliary chain frozen at parameter $\rho_{k-\tau_{k}}$; see also the proof of (Zeng et al., 2024, Lemma 1). Specifically, define